OpenStudy (anonymous):

A fair coin is tossed 3 times in a row. If X is the number of heads counted, what is the mean of the probability distribution of X? 1.0 1.5 2.0 2.5 3.0

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please wait a moment

OpenStudy (michele_laino):

please wait, your last equation is the mean value of x, so it can not be equal to 1

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

I'm trying to solve your question, please wait...

OpenStudy (anonymous):

Yea its all good

OpenStudy (michele_laino):

X can be 0, 1, 2, or 3

OpenStudy (michele_laino):

so we get the corresponding probabilities: X=0--->p=0 X=1--->p=1/3 X=2--->p=2/3 X=3--->p=1

OpenStudy (anonymous):

so it is 1.0?

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

cause x=3 at p=1

OpenStudy (michele_laino):

we have to determine the mean value of X

OpenStudy (anonymous):

so find the mean?

OpenStudy (anonymous):

I get a 0.5 for the mean

OpenStudy (michele_laino):

yes! I got the same result

OpenStudy (anonymous):

maybe the answer is 1

OpenStudy (michele_laino):

I think that we have to consider the binomial distribution.

OpenStudy (michele_laino):

now, if I toss my coin three times in a row, then we have N tests, and X is the number of success in our tests, furthermore the probability to get a tail is 1/2, and the probability to get a head is 1/2, so the probability to get X head is given by the binomial distribution, as follows: \[\left( {\begin{array}{*{20}{c}} 3 \\ X \end{array}} \right){\left( {\frac{1}{2}} \right)^3}\]

OpenStudy (michele_laino):

where: \[\left( {\begin{array}{*{20}{c}} 3 \\ X \end{array}} \right) = \frac{{3!}}{{X!\left( {3 - X} \right)!}}\]

OpenStudy (anonymous):

I get (6 (3-X)!)/(X!)

OpenStudy (michele_laino):

no, please, the mean value is given by the subsequent formula: \[\mu = \sqrt {Npq} = \sqrt {3 \times \frac{1}{2} \times \frac{1}{2}} = ...?\]

OpenStudy (anonymous):

0.866

OpenStudy (anonymous):

or sqrt(3)/2

OpenStudy (michele_laino):

Sorry I have made an error, the mean value of the binomial didtribution, is given by the subsequent formula: \[\Large \mu = Np = 3 \times \frac{1}{2} = ...?\]

OpenStudy (michele_laino):

distribution*

OpenStudy (anonymous):

1.5

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

p is the probability to get a head, namely p=1/2, whereas q is the probability to get a tail, namely q=p=1/2

OpenStudy (anonymous):

I also have more questions but I don't think you would want to help me again

OpenStudy (anonymous):

thanks alot

OpenStudy (michele_laino):

I have to help another student, so you have to wait, please post your question

OpenStudy (michele_laino):

ok! I can help you now!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What is the mean of this probability distribution?

OpenStudy (anonymous):

P(X=xi) is for the decimals :0.2,0.3,0.2,0.1,0.1,0.1

OpenStudy (anonymous):

1.9 2.9 3.1 3.2 3.4 are the answers

OpenStudy (michele_laino):

here we have to apply the classical definition of mean value, so we have to do this computation: \[\large 1 \times 0.2 + 2 \times 0.3 + 3 \times 0.2 + 4 \times 0.1 + 5 \times 0.1 + 6 \times 0.1 = ...?\]

OpenStudy (anonymous):

2.9

OpenStudy (anonymous):

1*0.2+2*0.3+3*0.2+4*0.1+5*0.1+6*0.1=2.9

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

one last question

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

I think uniform means continuous

OpenStudy (michele_laino):

I think that your distribution is like this: 

OpenStudy (anonymous):

yea

OpenStudy (michele_laino):

namely: p(x)=0, if x>b or x<a, and p(x)= b-a, if a<x<b

OpenStudy (anonymous):

we can eliminate a

OpenStudy (michele_laino):

oops. I have made an error, here is the right graph: